OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6). 7 -14 0 14

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

slope = y2 - y1 / x2 - x1

OpenStudy (anonymous):

would be 7 then

OpenStudy (anonymous):

so that would give (-6 - 8) / (-3-4) = ?

OpenStudy (anonymous):

i got 7

OpenStudy (anonymous):

now that would give you 0

OpenStudy (anonymous):

sorry, it should be (8 -(-6))/ (-4-(-3))=

OpenStudy (anonymous):

how could 14/7 be 0?

OpenStudy (anonymous):

oh it =2

OpenStudy (anonymous):

if you work out the negative signs you'll get (8+6)/(-4+3) = ?

OpenStudy (anonymous):

oh i did the opposite

OpenStudy (anonymous):

yeah, it's 2, but 2 is not an option, which means that the first point mentioned is actually the second point in the formula

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

(-4,8)= (x2, y2)

OpenStudy (anonymous):

(-3,-6)=(x1, y1) if the slope formula = (y2 - y1) / (x2 - x1) then you just plug in the numbers, and simplify the negative signs... + times - = - - times - = + + times + = +

OpenStudy (anonymous):

i got 0

OpenStudy (anonymous):

you can't have 0 as a result from a division, unless the nominator is 0

OpenStudy (anonymous):

hum

OpenStudy (anonymous):

\[slope = \frac{ y _{2} - y _{1} }{ x _{2}-x _{1} }\] \[slope = \frac{ 8 - (-6) }{ -4-(-3) }\] \[slope = \frac{ 8 + 6 }{ -4+ 3 }\] you first add up what is in the numenator, secondly you add up what is in the denominator, and then you divided those 2 results which is?

OpenStudy (anonymous):

so it 14 thats what i got

OpenStudy (anonymous):

14/(-1) = -14 ;-)

OpenStudy (anonymous):

-4 + 3 = -1 when you divide by -1 it's like dividing by 1, but you change the sign from positive to -